World Auto uses the formula p + -3000n + 17,000 to determine the wholesale price for a used Ford Focus, where P is the price in dollars and n is the age of the car in years.
a.) find the price for a Focus that is 1,2, or 3 years old
b.) Find the slpoe and p- intercept for the line
P = -3000n+ 17,000.
c.) Interpret the slope and P -intercept
0 solutions

Answer by Mathtut(3670)

a)just plug in 1,2 and 3 for n
:
I will do the 2nd one(2years old)
:
P=-3000(2)+17000=$11000 wholesale price after 2 years
:
b)remember this is written in slope intercept form y=mx+b where m is the slope and b is the y or in ur case the p intercept
:
so slope is -3000 and the p intercept is 17000
:
c)the p intecept represents the beginning value of the Focus at year 0
the slope represents the change in value each year. In other words each year the Focus loses $3000 worth of wholesale value.